OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. Type your answers in the blanks. Show your work.|4x+3|=9+2x x = _____ or _____

OpenStudy (anonymous):

@amberxoxo

OpenStudy (anonymous):

@KamiBug

OpenStudy (freckles):

you have two equations to consider 4x+3=9+2x ; 4x+3=-(9+2x) solve both of these and then check the outcomes

OpenStudy (anonymous):

my brain is going to exlode first day back and i have to do all this but my techer hasnt reallly tuaght us jank about it

OpenStudy (anonymous):

any idea @amberxoxo

OpenStudy (anonymous):

lol hold on

OpenStudy (anonymous):

ughh sh*t idk man

OpenStudy (anonymous):

lol same here

OpenStudy (anonymous):

@ganeshie8 @girlover

OpenStudy (freckles):

do you not want to attempt to solve linear equations above?

OpenStudy (freckles):

4x+3=9+2x ; 4x+3=-(9+2x) 4x+3=-9-2x <--distributed here Get your terms with the x on one side and get your terms without x on the opposite side

OpenStudy (freckles):

by addition or subtraction

OpenStudy (anonymous):

naw i think you got this partner lol plz help me i will try my best to do some of the work but i really dont know what im doing

OpenStudy (anonymous):

by addition right

OpenStudy (freckles):

For 4x+3=9+2x Subtract 2x on both sides Then subtract 3 on both sides can you do that?

OpenStudy (anonymous):

i would like to but i dont wanna mess it up plus i hate to mess up your flow waiting on me your doing such a really good job

OpenStudy (freckles):

I will probably help someone else who wants to try.

OpenStudy (freckles):

Good luck.

OpenStudy (anonymous):

ok @KamiBug

OpenStudy (anonymous):

thnx for the luck

OpenStudy (anonymous):

@justnick09

OpenStudy (anonymous):

this is some bs

OpenStudy (anonymous):

lol hell yes

OpenStudy (anonymous):

@BabyGirlAj @ashhhrodriguez @willsgirl16 @ilovewolf

OpenStudy (anonymous):

@micahm

OpenStudy (anonymous):

@morganchavez